OpenStudy (loser66):

what's wrong ? \[4x^2-4x-3=0\]

OpenStudy (loser66):

\[(2x)^2-4x-3=0\] (2x-1)^2=4 2x-1=2 x=3/2

OpenStudy (loser66):

I

OpenStudy (loser66):

miss x =-1/2 but wha

OpenStudy (rsadhvika):

looks okay ?

OpenStudy (loser66):

'*what's wrong

OpenStudy (loser66):

bad computer, it breaks my sentence, sorry for that

OpenStudy (rsadhvika):

looks correct. i see nothing wrong as such o.o

OpenStudy (loser66):

but if you use discriminant, you have other solution x= -1/2. by this way, it disappears

OpenStudy (cwrw238):

you missed the negative square root

OpenStudy (loser66):

yes. thank you

OpenStudy (cwrw238):

yw

OpenStudy (rsadhvika):

you need to sleep more :)

OpenStudy (cwrw238):

lol

OpenStudy (loser66):

hey, you didn't find it out as me. who needs sleep

OpenStudy (rsadhvika):

you misunderstood me, i interpreted u got two solutions already, and you have something more to say

OpenStudy (loser66):

ok, thanks you guys. I have a bunch of homework to do, cannot sleep

OpenStudy (rsadhvika):

:D